### v1

# EmoLoom-2B: Fast Base-Model Screening for Emotion Classification and VAD with Lexicon-Weak Supervision and KV-Off Evaluation

Abstract
We present EmoLoom-2B, a lightweight and reproducible pipeline that turns small language models ($\sim$2B) into fast-to-screen candidates for joint emotion classification and Valence–Arousal–Dominance (VAD). To eliminate protocol drift, we unify training and inference with a one-line JSON I/O contract and report a ParseOK rate alongside standard task metrics. Fairness is further enforced by a public KV-off decoding setting shared across training and evaluation. Around this backbone, we add two orthogonal regularizers: a VAD-preserving constraint that aligns the VAD implied by the generated text with the target triplet, and a lightweight external appraisal-atom verifier (goal attainment, controllability, certainty, fairness) that guides training without lengthening justifications. We also introduce a simple Valence Flip augmentation and an A/B mixture schedule with entropy-aware temperature cooling to trade off coverage and convergence during SFT. Using Qwen-1.8B-Chat as the base model, we train on GoEmotions and EmpatheticDialogues and probe cross-corpus generalization on DailyDialog with lexicon-derived weak VAD. On development sets, EmoLoom-2B attains Macro-F1 of 0.35 and VAD $(1-\mathrm{RMSE})$ of 0.94 with ParseOK =1.00=1.00; a time-bounded cross-corpus quick evaluation yields Macro-F1 of 0.31 while maintaining robust validity. The recipe is budget-aware, auditable, and re-entrant, providing a dependable screening pass before heavier training or multimodal fusion.
Preprint.
Introduction
Understanding human emotion from language remains a central capability for socially aware AI systems, with practical impact on mental-health support, education, safety moderation, and affective conversational agents. In low-latency or resource-constrained settings, small language models (SLMs, $\sim$2B parameters) are especially attractive. Yet, despite rapid progress, the evaluation of emotion understanding—especially when classification must co-exist with continuous Valence–Arousal–Dominance (VAD) regression—often suffers from three persistent issues: (i) protocol drift between training and inference that turns formatting and parsing into confounders rather than signal; (ii) fairness gaps in decoding (e.g., inconsistent use of KV-cache and generation hyperparameters) that inflate or deflate scores; and (iii) weak-supervision pipelines that do not explicitly preserve VAD semantics in the generated answers, undermining consistency even when label accuracy looks strong.
We introduce EmoLoom-2B, a lightweight and reproducible pipeline that turns a broad pool of open SLMs into fast-to-screen candidates for emotion classification and VAD, while enforcing a protocol-true and fair evaluation. The core idea is to unify data, training, and inference under a one-line JSON I/O contract and to remove avoidable evaluation variance by adopting KV-off decoding as a default public setting. Around this spine, we incorporate two orthogonal semantic regularizers. First, a VAD-preserving constraint aligns the VAD implied by the generated text with the target VAD triplet, stabilizing continuous predictions without injecting long rationales. Second, a lightweight appraisal-atom verifier—instantiated as a compact external classifier over goal attainment, controllability, certainty, fairness, and related “cognitive appraisal” factors—provides training-time guidance without entangling the generator with lengthy natural-language justifications. To improve polarity sensitivity, we introduce a simple Valence Flip augmentation that creates polarity-mirrored pairs and encourages symmetric responses. Finally, to balance coverage and convergence during SFT, we use A/B mixture sampling with entropy-aware temperature scheduling: high-entropy samples dominate early to widen coverage; the schedule cools over training to consolidate reliable patterns. An overview is shown in Fig. 1.
Our practical motivation is twofold. First, many downstream deployments require a dependable screening pass to select a base model before heavier training or multi-modal fusion. Such screening should be budget-aware, transparent, and stable across machines. Second, when emotion categories and VAD must be produced jointly, the system should optimize for both accuracy and answer validity: the output must be parseable, internally consistent, and semantically faithful to the requested format. EmoLoom-2B operationalizes these requirements through a minimal contract: one-line JSON with fields for multi-label labels, continuous vad={v,a,d}, and a short rationale. We report standard metrics (Macro-F1/Precision/Recall for multi-label classification; $1-\mathrm{RMSE}$ for VAD) together with a ParseOK rate that quantifies adherence to the output contract.
Concretely, we systematically screen candidate SLM backbones and adopt Qwen-1.8B-Chat as the default base due to its favorable coverage–stability trade-off in our setup. Training uses common, steady-state practices (bf16/TF32, gradient checkpointing, cache-robust allocation) and the exact decoding configuration used for evaluation (use_cache=false), eliminating train–test discrepancies. Datasets cover GoEmotions and EmpatheticDialogues for training and development, complemented by a cross-corpus probe on DailyDialog with lexicon-driven weak VAD to assess generalization under domain shift. All scripts follow a single manifest and directory layout to make runs auditable and re-entrant.
Contributions.
This work offers a compact recipe for dependable, budget-aware emotion modeling with SLMs:
  1. 1.
    Protocol-true I/O. A one-line JSON contract unifies training and inference. We pair standard task metrics with ParseOK to directly measure contract adherence, reducing “format wins” that do not reflect genuine modeling gains.
  2. 2.
    Fair public decoding. We institutionalize KV-off decoding as the default evaluation mode and match training/inference decoding settings, removing a common source of variance and improving replicability across machines and seeds.
  3. 3.
    VAD-preserving regularization. A simple consistency loss aligns the VAD implied by the generated answer with the target VAD, improving continuous estimates without relying on lengthy, hard-to-grade rationales.
  4. 4.
    Lightweight appraisal verifier. A compact, external classifier over cognitive appraisal “atoms” provides training-time guidance, decoupling semantic checks from the generator and avoiding justification length as a hidden knob.
  5. 5.
    Polarity symmetry via Valence Flip. Polarity-mirrored pairs encourage sensitivity to valence and yield a diagnostic handle on polarity robustness at evaluation time.
  6. 6.
    Coverage–convergence scheduling. A/B mixture sampling with an entropy-aware temperature schedule stabilizes small-model SFT by emphasizing diverse, informative samples early and consolidating later.
  7. 7.
    Fast screening under time budget. A clear, time-bounded evaluation path enables quick comparison of candidate backbones before investing in heavier training or multi-modal integration, with unified logs, figures, and tables for paper-ready reporting.
Empirical preview.
Across development sets and a cross-corpus probe, EmoLoom-2B improves Macro-F1 and VAD error while maintaining high ParseOK; qualitatively, we see fewer format failures and steadier valence under polarity stress-tests. A 20:80 mix often gives the best F1–VAD trade-off, though dataset-dependent; detailed results and ablations follow.
Scope and ethics.
We focus on text-based emotion understanding in English. Weak labels derived from public lexica are used only for training signals and diagnostics; we release aggregated metrics, not raw personal data. The pipeline is designed for reproducibility and auditability, with identical decoding for training and evaluation.
Paper roadmap.
Related Work reviews prior art; Method covers the I/O contract, VAD-preserving loss, appraisal verifier, Valence Flip, and the schedule (Fig. 1). Data and Weak-Label Generation details datasets and weak-label conversion; Implementation Details presents screening. Evaluation Protocol defines metrics; Results, Ablations and Sensitivity, and Discussion and Limitations report findings and analyze design choices; Conclusion closes.
Figure 1: EmoLoom-2B overview. Left: Existing issues. When jointly asking for emotion labels and VAD, big models often drift in protocol, rely on subjective decoding, or fail to preserve VAD semantics. Right: Workflow (A–D). A Model selection among open SLMs; B Zeroshot sanity checks on target datasets; C Training on Qwen-1.8B-Chat with mixed-extract splits (20:80/50:50/80:20) and CE+VAD loss under protocol-true JSON I/O; D Evaluation with Macro-F1/P/R, VAD ($1-\mathrm{RMSE}$), cross-corpus probe, and ParseOK validity. We adopt KV-off decoding for fairness across training and evaluation.
Related Work
LLM-based emotion classification and VAD.
Early neural approaches treated emotion as single- or multi-label classification, later extending to joint prediction of discrete categories and continuous Valence–Arousal–Dominance (VAD). Recent small language models (SLMs, $\sim$2B) inherit strong text priors and can be instruction-tuned for the joint task, but evaluation often confounds modeling quality with formatting or decoding choices. Common resources include crowd-labeled dialogue and sentence datasets (e.g., GoEmotions, EmpatheticDialogues, DailyDialog) and have encouraged multi-label metrics (Macro-F1/Precision/Recall) alongside continuous VAD error. Our work follows this joint formulation but emphasizes answer validity via a one-line JSON contract and a ParseOK rate so that improvements are not artifacts of prompt or parser idiosyncrasies (dg:20; rs:19; ls:17; r:80; m:18; wkb:13).
Appraisal Theory in computational emotion.
Psychological theories describe emotions through cognitive appraisals such as goal attainment, controllability, certainty, and fairness. Computational uses range from feature engineering and rule systems to text classifiers trained on appraisal cues. Many LLM pipelines now generate lengthy rationales to imitate human explanations, yet rationale length becomes a tuning knob and may not translate to better VAD fidelity. We instead adopt appraisal theory as a training-time constraint: a lightweight external “appraisal-atom” verifier provides semantic guidance without fusing explanations into the generator, preserving simplicity, speed, and controllable failure modes (occ:88; se:85; scherer:01; e:92).
Lexicon-based weak supervision and VAD norms.
Emotion and affective lexica (e.g., NRC Emotion Lexicon) and affective norm lists (e.g., Warriner’s VAD norms) have long enabled low-cost labels or priors. They are effective for coverage and quick diagnostics but are sensitive to domain shift, polysemy, and context (negation, sarcasm). In EmoLoom-2B, lexica are not treated as fixed labels; they supply weak VAD signals to (i) regularize the model through a VAD-preserving consistency loss that aligns generated text with target VAD and (ii) support polarity stress tests via Valence Flip augmentation. This keeps lexical priors useful while avoiding overcommitment to word-level heuristics (mt:13; wkb:13; m:18; pz:06; wb:10; pl:08; rb:20).
Small-model SFT and evaluation fairness.
Instruction/SFT pipelines for small models typically vary decoding hyperparameters (temperature, top-$p$, max length) and cache usage across training and evaluation, yielding non-trivial score variance and limited replicability. KV-cache in particular changes compute and sometimes output trajectories across hardware and seeds, complicating fair comparison. We advocate a protocol-true setup: identical JSON I/O and KV-off decoding for both training and evaluation, making results comparable across machines and allowing ablations to measure modeling changes rather than decoding drift (bm:20; dc:19; dy:19; hbd:20; liang:22; ba:21; p:21; bgm:21).
Quick, budget-aware evaluation.
Many deployments need a rapid “screening pass” to choose a base model before heavier training or multimodal fusion. Prior work on efficient benchmarking suggests time- or compute-bounded evaluation to trade breadth for speed, but such protocols are rarely standardized for emotion + VAD. EmoLoom-2B provides a simple, auditable quick-eval path—shared metrics, unified logs/figures, and ETA reporting—so candidate backbones can be compared under the same budget (liang:22; rwg:20).
Positioning.
Unlike frameworks that replicate full psychological pipelines or rely on explanation-conditioned generation, we use psychology as structured constraints on the training objective: a VAD-preserving loss and a compact appraisal verifier. Together with a protocol-true JSON contract and KV-off decoding, this yields a small, fast, and reproducible recipe for joint emotion/VAD modeling that is easy to audit and extend.
Method
Task and One-line JSON I/O
We jointly predict multi-label emotions and continuous Valence–Arousal–Dominance (VAD) per utterance $x$. The model must return a single-line JSON. To avoid overfull boxes in two-column layout, we typeset it with automatic line breaking:
Listing 1: One-line JSON I/O (wrapped for typesetting)
{"labels":["disgust"],"vad":{"v":0.42,"a":0.21,"d":0.49},"rationale":"..."}
where $\mathrm{vad}\in[0,1]^{3}$ is rounded to two decimals and rationale is a short English phrase. During evaluation we tail-scan the generated text to find a valid JSON object; success defines ParseOK. Formally, with $N$ samples,
$\mathrm{ParseOK}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\left\{\mathrm{parse}(\hat{y}_{i})\right\}.$
We compute task metrics on valid outputs only and report ParseOK side-by-side to expose formatting failures. Training and inference use the same prompt and decoding; KV cache is disabled for fairness (Sec. Fair KV-off Protocol and Reproducibility).
Label and VAD losses.
Let $K$ be the number of emotion labels, $y\in\{0,1\}^{K}$ the multi-hot target, and $p\in(0,1)^{K}$ the predicted probabilities. We use a standard multi-label BCE:
$\mathcal{L}_{\mathrm{cls}}=\frac{1}{K}\sum_{k=1}^{K}\Big[-y_{k}\log p_{k}-(1-y_{k})\log(1-p_{k})\Big].$
For VAD regression with target $\hat{\boldsymbol{v}}\in[0,1]^{3}$ and prediction $\boldsymbol{v}\in[0,1]^{3}$,
$\mathcal{L}_{\mathrm{reg}}=\|\boldsymbol{v}-\hat{\boldsymbol{v}}\|_{2}^{2}.$
VAD-Preserving Consistency
To encourage semantic faithfulness, we align the VAD implied by the generated text with the numeric VAD. Let the output text be $a$, tokenized to $\{t_{j}\}$; given a lexicon $\ell(t)$ that maps tokens to VAD scores (missing tokens ignored), we aggregate
$\boldsymbol{v}_{\text{text}}(a)=\frac{\sum_{j}\omega_{j}\,\ell(t_{j})}{\sum_{j}\omega_{j}},\quad\omega_{j}=\mathbf{1}\!\left\{\,\ell(t_{j})\ \mathrm{exists}\,\right\},$
and define the consistency loss
$\mathcal{L}_{\mathrm{vad}}=\|\boldsymbol{v}_{\text{text}}(a)-\hat{\boldsymbol{v}}\|_{2}.$
Lightweight Appraisal-Atom Verifier
Inspired by Appraisal Theory, we build a compact external verifier over $M$ appraisal “atoms” (goal attainment, controllability, certainty, fairness, etc.). Let $\boldsymbol{s}\in[0,1]^{M}$ be the verifier scores on $(x,a)$ and $\tilde{\boldsymbol{s}}(y)$ the atom prototype derived from the gold labels (or weak rules). We penalize mismatches via a logistic loss:
$\mathcal{L}_{\mathrm{app}}=\frac{1}{M}\sum_{m=1}^{M}\Big[-\tilde{s}_{m}\log s_{m}-(1-\tilde{s}_{m})\log(1-s_{m})\Big].$
The verifier is trained separately as a LogReg/MLP over weak features and used only as a training-time constraint; it is not concatenated to the generator output, avoiding explanation-length as a hidden knob.
Valence Flip Symmetry
[…]
$\mathcal{L}_{\mathrm{flip}}=\big|\,\big(v(x)-\tfrac{1}{2}\big)+\big(v(x^{\prime})-\tfrac{1}{2}\big)\,\big|.$
This acts as a small regularizer and a diagnostic stress test of polarity sensitivity. (Only valence is constrained; arousal/dominance remain free.)
A/B Mixture with Entropy-Aware Temperature
We study three mixture ratios between GoEmotions and EmpatheticDialogues: 20:80, 50:50, 80:20. At each step we choose the next sample source $s\in\{A,B\}$ via
$p(s)=\mathrm{softmax}\!\left(\frac{w_{s}/\mathrm{conf}_{s}}{T}\right),$
where $w_{s}$ is the target ratio weight, $\mathrm{conf}_{s}$ is a running confidence proxy (e.g., moving-average entropy), and $T$ cools linearly across training: $T_{t}=T_{0}-(T_{0}-T_{1})\frac{t}{T_{\max}}$. This schedule emphasizes broad coverage early and consolidation later. Empirically, the 20:80 configuration yields the best F1–VAD trade-off (Macro-F1 0.3500, $1\!-\!\mathrm{RMSE}_{\text{VAD}}=0.9417$, ParseOK=1.000), with quick cross-corpus performance $\approx\!0.31$ Macro-F1 on DailyDialog under a 1h budget.
Fair KV-off Protocol and Reproducibility
To remove decoding-induced variance, we disable the key–value cache and keep decoding identical for training and evaluation (use_cache=false). We also enable gradient checkpointing and an OOM self-healing routine that decreases max_len and increases grad_accum on failure, then resumes training. These settings are part of the public evaluation recipe and were used for all reported numbers.
Determinism and versioning. We fix RNG seeds for Python/NumPy/PyTorch ({11, 22, 33}), enable cudnn.deterministic=True and disable cudnn.benchmark. Each checkpoint bundles the exact config (YAML/JSON), tokenizer hash, and git commit.
Environment stamp. Runs log GPU model, driver, CUDA/cuDNN, and CPU; the eval script asserts major-version compatibility and warns on drift to avoid accidental speed/quality changes.
Data integrity. All shards carry SHA-1 checksums and a frozen permutation index for train/dev; the loader asserts counts and de-duplicates by sha1(text||id) so that re-sharding cannot change splits.
Decoding invariants. We use deterministic greedy decoding (temperature=0, top_p=1.0, EOS early stop) with KV-off; identical prompt templates and a single-line JSON schema ensure protocol-true comparison across ablations.
[…]
Precision and numerics. Training uses bf16 forward with TF32 matmul on Ampere/ADA; gradients are clipped at $1.0$ and optimized by AdamW ($\beta_{1}{=}0.9,\ \beta_{2}{=}0.95$, weight decay $0.1$). The LR follows cosine decay with $3\%$ warmup; we checkpoint EMA weights for evaluation parity.
Prompt/length invariants. A single frozen prompt template is used across all runs; generation uses a fixed budget of $64$ tokens with early stop on } or EOS. Inputs are truncated to max_len=1536=1536 tokens (prompt+context), and any overflow is counted in the ParseOK denominator.
Overall objective.
The final loss combines all components,
$\mathcal{L}=\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}+\lambda_{\mathrm{reg}}\mathcal{L}_{\mathrm{reg}}+\lambda_{\mathrm{vad}}\mathcal{L}_{\mathrm{vad}}+\lambda_{\mathrm{app}}\mathcal{L}_{\mathrm{app}}+\lambda_{\mathrm{flip}}\mathcal{L}_{\mathrm{flip}},$
with $\lambda$’s selected on the dev set.
Algorithm 1 Protocol-true SFT with KV-off and OOM self-healing
1:Datasets $\mathcal{D}_{A},\mathcal{D}_{B}$ (GoEmo/Empathetic), weights $w_{A},w_{B}$; NRC-VAD lexicon $\ell(\cdot)$; verifier $f_{\text{app}}$; loss weights $\lambda_{\text{cls}},\lambda_{\text{reg}},\lambda_{\text{vad}},\lambda_{\text{app}},\lambda_{\text{flip}}$; schedule $T_{0}\!\to\!T_{1}$; seed
2:Trained parameters $\theta$ (Qwen-1.8B-Chat), JSON-valid decoding
3:Initialize $\theta$; set use_cache=false; enable grad checkpointing; set max_len, grad_accum
[…]
5:for $t=1$ to $T_{\max}$ do
6:  $T\leftarrow T_{0}-(T_{0}-T_{1})\,t/T_{\max}$ $\triangleright$ linear temperature cooling
7:  $s\sim\mathrm{softmax}(\tfrac{w_{s}/\text{conf}_{s}}{T})$, $s\in\{A,B\}$
8:  $(x,y,\hat{\boldsymbol{v}})\sim\mathcal{D}_{s}$; optionally flipped $(x^{\prime},y^{\prime},\hat{\boldsymbol{v}}^{\prime})$
9:  $a,\boldsymbol{v}\leftarrow\text{Decode}_{\text{KV-off}}(x;\theta)$
10:  $\boldsymbol{v}_{\text{text}}\leftarrow\frac{\sum_{j}\omega_{j}\,\ell(t_{j})}{\sum_{j}\omega_{j}}$ from tokens $\{t_{j}\}$ in $a$
11:  $\mathcal{L}_{\text{cls}}\leftarrow\frac{1}{K}\sum_{k}[-y_{k}\log p_{k}-(1-y_{k})\log(1-p_{k})]$
12:  $\mathcal{L}_{\text{reg}}\leftarrow\|\boldsymbol{v}-\hat{\boldsymbol{v}}\|_{2}^{2}$; $\mathcal{L}_{\text{vad}}\leftarrow\|\boldsymbol{v}_{\text{text}}-\hat{\boldsymbol{v}}\|_{2}$
13:  $\mathcal{L}_{\text{app}}\leftarrow\frac{1}{M}\sum_{m}\text{BCE}(f_{\text{app}}^{(m)}(x,a),\tilde{s}_{m}(y))$
14:  $\mathcal{L}_{\text{flip}}\leftarrow|(v(x)-\tfrac{1}{2})+(v(x^{\prime})-\tfrac{1}{2})|$ (if flip exists)
15:  $\mathcal{L}\leftarrow\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}+\lambda_{\text{vad}}\mathcal{L}_{\text{vad}}+\lambda_{\text{app}}\mathcal{L}_{\text{app}}+\lambda_{\text{flip}}\mathcal{L}_{\text{flip}}$
[…]
17:  Update $\text{conf}_{s}\leftarrow\alpha\,\text{conf}_{s}+(1-\alpha)\exp(-H_{t})$
Algorithm 2 Advanced: Entropy-aware A/B mixing with online uncertainty and budget-aware early stopping
1:Dev sets $\mathcal{D}^{\text{dev}}_{A,B}$; budget $B$ (mins); windows $W_{\text{score}},W_{\text{eta}}$; threshold $\delta$; parser TailScanJSON
2:Best checkpoint $\theta^{\star}$ under budget, with metrics and ParseOK
[…]
4:for epoch $1,2,\ldots$ do
5:  Train one pass using Alg. 1; log entropy $\bar{H}_{s}$ and ParseOK $\pi$
[…]
7:  Eval dev with use_cache=false: decode $\to$ TailScanJSON; accumulate Macro-F1/P/R, $1-\mathrm{RMSE}_{\text{VAD}}$, $\pi$
8:  $\mathcal{S}\leftarrow\text{MacroF1}+\beta(1-\mathrm{RMSE})+\gamma\,\pi$
9:  keep sliding windows to get $\overline{\mathcal{S}}$ and median ETA
10:  if $\mathcal{S}>\mathcal{S}^{\star}$ then
11:   $\theta^{\star}\leftarrow\theta$; $\mathcal{S}^{\star}\leftarrow\mathcal{S}$
12:  else if $\overline{\mathcal{S}}<\mathcal{S}^{\star}+\delta$ and median(ETA)>> remaining budget then
13:   break $\triangleright$ budget-aware early stop
14:  Rebalance mixing: $w_{s}\leftarrow\mathrm{Normalize}\{w_{s}/(\bar{H}_{s}+\epsilon)\}$
15:return $\theta^{\star}$ with final metrics and ParseOK
Data and Weak-Label Generation
Training/Dev.
We use GoEmotions and EmpatheticDialogues. Utterances are filtered by quality-control flags (qc_flags), length $[3,128]$ tokens, and deduplicated via sha1(text||id). We compute a lexicon-coverage confidence
$\mathrm{vad\_conf}(x)=\frac{1}{J}\sum_{j=1}^{J}\mathbf{1}\!\left\{\,\ell(t_{j})\ \mathrm{exists}\,\right\},$
and keep samples with $\mathrm{vad\_conf}\geq\tau$ (default $\tau\in\{0.75,0.80\}$). Datasets are then mixed by ratios 20:80 / 50:50 / 80:20 with a fixed split ($\mathrm{dev\_frac}\approx 5\%$) and seed for exact reproducibility. The resulting dev-set sizes used in Sec. Results are:
  • 20:80 $\Rightarrow n_{\mathrm{dev}}=3663$,
  • 50:50 $\Rightarrow n_{\mathrm{dev}}=3309$,
  • 80:20 $\Rightarrow n_{\mathrm{dev}}=2068$.
We report Macro-F1/Precision/Recall, VAD $(1-\mathrm{RMSE})$, and ParseOK. The best overall configuration is 20:80 (dg:20; rs:19).
Weak VAD normalization.
For tokens with NRC-VAD entries, we use the provided $[0,1]$ scores; for lexica in $[1,9]$ (e.g., Warriner) we min–max normalize as $\tilde{v}=(v-1)/8$ and clip to $[0,1]$, then apply a small label-smoothing $\epsilon{=}0.01$:
$\hat{\boldsymbol{v}}\leftarrow(1-2\epsilon)\,\hat{\boldsymbol{v}}+\epsilon.$
Utterance-level weak VAD is aggregated by weighted mean over covered tokens (weights $\omega_{j}{=}1$); samples with zero coverage are excluded by the $\mathrm{vad\_conf}$ filter (m:18; wkb:13).
Cross-corpus probe (weak VAD).
For DailyDialog, we split dialogues into utterances, compute weak VAD via the same aggregation, and map its emotion tags into our label space (unseen tags $\rightarrow$ other). The converted file is exported under a distinct name (e.g., dev_dailydialog_weak.jsonl) to avoid overwriting the main dev set. Under a 1-hour quick-eval regime, the 20:80 model reaches Macro-F1 $\approx 0.307$ and ParseOK $\approx 0.976$ (ls:17).
File layout and manifests.
We keep a flat, auditable directory with frozen names; the manifest lists all shards and their checksums:
data/
goemotions/train.jsonl
goemotions/dev.jsonl
empathetic/train.jsonl
empathetic/dev.jsonl
dailydialog/dev_dailydialog_weak.jsonl
manifests/
mix_20_80.json mix_50_50.json mix_80_20.json
checksums.sha1
Each manifest stores the exact paths, mixture ratio, seed, and split sizes:
{"mix":"20:80","seed":42,"dev_frac":0.05,
"sources":{"A":"data/goemotions/train.jsonl",
"B":"data/empathetic/train.jsonl"},
"dev":{"A":"data/goemotions/dev.jsonl",
"B":"data/empathetic/dev.jsonl"}}
Schema of a JSONL record.
Each line carries gold multi-labels, weak VAD, and optional QC flags:
{"id":"goe_001122","text":"I am not happy about this.",
"labels":["disgust","anger"], "vad":{"v":0.21,"a":0.58,"d":0.43},
"qc_flags":{"lang":"en","len":6,"dedup":false}, "split":"train"}
During training/evaluation, the model must emit a single-line JSON adhering to the schema in Listing 1 (Sec. Task and One-line JSON I/O); validity is counted by ParseOK.
Implementation snapshot.
We select Qwen-1.8B-Chat as the sole backbone after screening and keep reproducible defaults (bf16/TF32, gradient clipping $1.0$, AdamW with cosine LR and $3\%$ warmup, gradient checkpointing, max length $1536$ with automatic downshift on OOM). All exported figures/tables are regenerated from JSON logs, and per-example predictions are released for auditability.
Implementation Details
Backbone Screening
We compared two 1–2B backbones, Qwen-1.8B-Chat and InternLM2-1.8B-SFT, under the protocol-true setup (JSON I/O + KV-off) and a quick-eval budget. A composite score was used to rank candidates:
$\mathrm{Score}=$  $\;0.4\,z(\mathrm{MacroF1})+0.4\,[\,z(\rho_{\mathrm{VAD}})-z(\mathrm{RMSE}_{\mathrm{VAD}})\,]$
$+0.2\,z(\mathrm{Quality})$
where $z(\cdot)$ denotes z-score across candidates. Figure 2 summarizes the normalized coverage; Table 1 lists the quick-eval numbers. Qwen shows stronger classification and VAD ( z=+1.0z{=}{+}1.0 ) while InternLM2 outputs slightly cleaner structure/JSON, yielding a lower composite. We therefore select Qwen-1.8B-Chat as the single backbone for all subsequent experiments.
[…]
Table 1: Backbone model selection (quick-eval). : mean across seed trio; : lower is better; §: normalized to $[0,1]$; : z-score across candidates.
model macro_f1 rmse_mean†‡ rho_mean quality§ q_json_ok q_struct_ok q_rat_len_ok $z_{\text{cls}}^{\star}$ $z_{\text{vad}}^{\star}$ $z_{\text{qual}}^{\star}$ composite
qwen-1.8b-chat 0.0403 0.2586 0.2407 0.3958 0.6221 0.2754 0.0107 +1.0 +1.0 1.0-1.0 +0.6
internlm2-1.8b-sft 0.0214 0.2747 0.2272 0.5024 0.7383 0.3564 0.1318 1.0-1.0 1.0-1.0 +1.0 0.6-0.6
Notes. Quick-eval uses KV-off, greedy decoding, and a fixed time budget; quality aggregates JSON parse, structural validity, and short-rationale preference.
Stability and Efficiency
Training uses bf16 forward with TF32 matmul; AdamW ($\beta_{1}{=}0.9,\ \beta_{2}{=}0.95$, weight decay 0.1), cosine LR with $3\%$ warmup, gradient clipping at 1.0, gradient checkpointing, and the OOM self-healing routine (Sec. Fair KV-off Protocol and Reproducibility). Figure 3 shows smoothed loss; the 20:80 mix converges quickest and to the lowest loss (min $\approx 0.160$), matching its best dev trade-off. Figure 4 visualizes LR schedule and gradient norms, indicating stable updates after the initial warmup (akin to curriculum-style scheduling (bengio:09)).
Figure 3: Training loss (smoothed). Curves correspond to mixtures 20:80 / 50:50 / 80:20. The 20:80 setting reaches the lowest minimum with the fastest decay.
Figure 4: LR & Grad-Norm panel. Top: cosine LR with $3\%$ warmup; Bottom: gradient norms stabilize quickly ($\sim$1.5–2.0) across mixtures, indicating healthy optimization under bf16+checkpointing.
Evaluation Protocol
Metrics and Gating
We evaluate on the subset of predictions that satisfy the JSON contract. Let the dataset size be $N$, the valid index set be $V$ with $|V|=N_{\!\text{val}}$. Task metrics are computed on $V$, while ParseOK is computed on all $N$.
ParseOK
$\mathrm{ParseOK}=\frac{N_{\!\text{val}}}{N}.$
Macro-F1 / P / R (gold subspace) Let $\mathcal{K}^{\ast}=\{k:\sum_{i=1}^{N}\mathbf{1}\!\left\{k\in Y_{i}\right\}>0\}$.
$\mathrm{TP}_{k}=\sum_{i\in V}\mathbf{1}\!\left\{k\in Y_{i}\cap\hat{Y}_{i}\right\}\,.$
$\mathrm{FP}_{k}=\sum_{i\in V}\mathbf{1}\!\left\{k\in\hat{Y}_{i}\setminus Y_{i}\right\}\,.$
$\mathrm{FN}_{k}=\sum_{i\in V}\mathbf{1}\!\left\{k\in Y_{i}\setminus\hat{Y}_{i}\right\}\,.$
$P_{k}=\frac{\mathrm{TP}_{k}}{\mathrm{TP}_{k}+\mathrm{FP}_{k}+\varepsilon}\,$,
$R_{k}=\frac{\mathrm{TP}_{k}}{\mathrm{TP}_{k}+\mathrm{FN}_{k}+\varepsilon}\,$,
$F1_{k}=\frac{2P_{k}R_{k}}{P_{k}+R_{k}+\varepsilon}\,$,
with $\varepsilon=10^{-9}$ for numerical stability.
$\mathrm{MacroP}=\frac{1}{|\mathcal{K}^{\ast}|}\sum_{k\in\mathcal{K}^{\ast}}P_{k}\,$,
$\mathrm{MacroR}=\frac{1}{|\mathcal{K}^{\ast}|}\sum_{k\in\mathcal{K}^{\ast}}R_{k}\,$,
$\mathrm{MacroF1}=\frac{1}{|\mathcal{K}^{\ast}|}\sum_{k\in\mathcal{K}^{\ast}}F1_{k}\,.$
VAD $(1-\mathrm{RMSE})$
$\mathrm{RMSE}_{\mathrm{VAD}}=\sqrt{\frac{1}{3N_{\!\text{val}}}\sum_{i\in V}\left\|\boldsymbol{v}_{i}-\hat{\boldsymbol{v}}_{i}\right\|_{2}^{2}}\,$,
$\mathrm{VAD}(1-\mathrm{RMSE})=1-\mathrm{RMSE}_{\mathrm{VAD}}\,.$
Two Evaluation Modes
[…]
Quick Eval (budgeted). Given a wall-clock budget $B$ minutes, evaluate a seeded, deterministic stream under the same decoding; print metrics online with ETA, and stop when time exceeds $B$. The final snapshot mirrors Full Eval fields but on $N_{\!\text{val}}^{(B)}\!\leq\!N_{\!\text{val}}$ and is used only for screening.
Reporting and Export
  • Contract-first logging: store commit, config hash, prompt template ID, and decoding config with metrics.
  • Unified artifacts: export .csv/.json for metrics and render figures/tables from the same sources.
  • Failure visibility: when $N_{\!\text{val}}<N$, also report label coverage and common parse failures.
Results
Training & Dev Summary. We summarize training dynamics (loss, LR, grad-norm; see Figs. 3, 4) and dev/cross-corpus results in Tables 2 and 3. Across mixtures, 20:80 achieves the lowest best loss and the strongest dev trade-off (Macro-F1 and VAD), under stable formatting in practice. Figures 56 visualize per-mixture trends.
Table 2: Training overview (trimmed). Lower is better for Best Loss. “Points” = #logged steps. best, second best.
EXP Points Best Loss Best Epoch
mix2080 90 0.1604 0.910
mix5050 82 0.1868 0.810
mix8020 46 0.2122 0.970
Table 3: Dev & cross-corpus evaluation (valid JSON only). best, second best.
EXP Macro-F1 Macro-R VAD(1-RMSE)
mix2080 0.3500 0.2693 0.9417
mix5050 0.3470 0.2657 0.9337
mix8020 0.3341 0.2509 0.9135
Quick-eval (DailyDialog, 1 h) for mix2080: Macro-F1 =0.3071=0.3071, VAD(1–RMSE) =0.8066=0.8066, n=6261n{=}6261.
Figure 5: Dev set scores (bars). mix2080 attains the top Macro-F1 and VAD.
Figure 6: Dev set radar. Metrics normalized to $[0,1]$ for shape comparison.
Ablations and Sensitivity
We ablate each component under the public KV-off protocol and the 20:80 mixture unless otherwise noted. Metrics are computed on valid JSON only; we observed ParseOK $\approx 1.00$ throughout.
(A) Remove VAD-preserving loss.
Dropping $\mathcal{L}_{\mathrm{vad}}$ increases VAD error and destabilizes polarity: $(1-\mathrm{RMSE})$ typically drops by $0.006\!\sim\!0.012$, Macro-F1 by $0.004\!\sim\!0.007$. Qualitatively we see more valence drift on polarity flips and slightly wider VAD spread.
(B) Remove appraisal-atom verifier.
Eliminating $\mathcal{L}_{\mathrm{app}}$ mainly hurts categories tied to controllability/fairness, with Macro-F1 falling by $0.010\!\sim\!0.015$ and negligible change in $(1-\mathrm{RMSE})$. Training wall-time is unaffected at inference (the verifier is not used at test).
(C) Remove Valence Flip.
Without flip pairs, polarity robustness degrades. Using paired samples $(x,x^{\prime})$, our symmetry diagnostic
$\mathcal{S}_{\mathrm{flip}}=\frac{1}{|\mathcal{P}|}\sum_{(x,x^{\prime})\in\mathcal{P}}\big|\,\big(v(x)-\tfrac{1}{2}\big)+\big(v(x^{\prime})-\tfrac{1}{2}\big)\,\big|$
rises from $\approx\!0.06$ to $\approx\!0.11$; Macro-F1 drops by $\sim\!0.003$.
(D) Temperature schedule.
Constant-$T$ under-covers early high-entropy regions (slower F1 gains). Early-cooling schedules overfit and lose $(1-\mathrm{RMSE})$ by $\sim\!0.005$. Linear cooling (our default) balances coverage and consolidation.
(E) Mixture ratio sensitivity.
Across 20:80, 50:50, 80:20, the 20:80 mixture yields the best Macro-F1 and $(1-\mathrm{RMSE})$ trade-off; differences between 20:80 and 50:50 are modest ($\leq 0.003$ F1; $\leq 0.008$ on $(1-\mathrm{RMSE})$).
(F) Loss weights.
Within a broad band ($\lambda_{\mathrm{vad}}\!\in[0.5,1.5]$, $\lambda_{\mathrm{app}}\!\in[0.5,1.0]$, $\lambda_{\mathrm{flip}}\!\in[0.2,0.6]$), performance is flat; overly large $\lambda_{\mathrm{vad}}$ can reduce label recall.
Discussion and Limitations
Weak-VAD bias. Lexicon-derived VAD is sensitive to domain shift, negation, and polysemy; our use is regularizing rather than prescriptive, and cross-corpus probes are reported separately.
Two-decimal quantization. Rounding VAD to two decimals improves contract adherence and reproducibility, but coarsens subtle affect. Future work can decouple the internal regressor from the printed precision.
Model capacity. With $\sim$2B parameters, nuanced appraisal interactions and long-range context remain challenging. Our recipe targets screening scenarios; scaling can combine this contract with larger backbones or multimodal inputs.
Protocol scope. The JSON I/O contract favors structured outputs; free-form empathy or long explanations are out-of-scope by design, though they can be layered on top in downstream systems.
Ethics and Societal Impact
We train on public datasets, release only aggregate metrics and anonymized IDs, and use weak VAD signals strictly for training-time regularization and diagnostics. The protocol avoids storing or redistributing personal texts; cross-corpus evaluation is inference-only. Potential risks include misinterpretation of affect in sensitive contexts and demographic skew in lexicons; we recommend deployment-time audits, opt-out mechanisms, and periodic re-evaluation on curated, balanced test suites.
Reproducibility and Checklist
Configs (YAML/JSON), tokenizer hashes, and git commits are bundled with checkpoints; seeds for Python/NumPy/PyTorch are fixed; decoding is deterministic (temperature=0, KV-off). Data manifests include SHA-1 checksums and frozen splits. One-click scripts reproduce the three mixtures, full/quick evaluations, and export figures/tables. The quick-eval path enforces a wall-clock budget and logs ETA for time-auditability.
Conclusion
A protocol-true JSON contract, a fair public KV-off decoding setup, and lightweight psychological constraints (VAD-preserving loss and an appraisal-atom verifier) provide a small, fast, and reproducible recipe for joint emotion classification and VAD with SLMs. The approach is budget-aware, easy to audit, and leaves headroom for multilingual and multimodal extensions.
Appendix A Appendix A: Implementation Details and Config
This appendix consolidates a no-source, minimal-yet-reproducible runbook. It specifies environment, data contract, CLI interfaces, and expected artifacts so an internal auditor can re-run our experiments with the private scripts that match these interfaces.
A.1 Hardware & OS. Single GPU with $\geq$24 GB VRAM; Ubuntu 20.04+; CUDA 12.x (driver matching).
A.2 Python Environment.
conda create -n emo_env python=3.12 -y && conda activate emo_env
pip install torch==2.3.1 --index-url https://download.pytorch.org/whl/cu121
pip install transformers==4.45.1 accelerate==0.34.2 datasets==2.20.0
pip install scikit-learn==1.4.2 matplotlib==3.8.4 pyyaml==6.0.1 tqdm==4.66.4
# Optional stability:
export PYTORCH_CUDA_ALLOC_CONF=expandable_segments:True
export CUBLAS_WORKSPACE_CONFIG=:16:8
export PYTHONHASHSEED=42
A.3 Project Layout.
/root/autodl-tmp/Emoloom-2B/
configs/ # YAML configs (hyperparameters)
data/
processed/
mix_20_80/ {train.jsonl, dev.jsonl}
mix_50_50/ {train.jsonl, dev.jsonl}
mix_80_20/ {train.jsonl, dev.jsonl}
raw/ dailydialog/ ... # raw corpora (optional for cross-corpus)
models/
qwen1_5_1_8b_chat/Qwen/Qwen1___5-1___8B-Chat/
outs/ runs/ # outputs and logs
dev.jsonl # DailyDialog converted for cross-corpus (optional)
src/ # private scripts (not disclosed)
A.4 Data Contract (JSONL). Each line contains at least:
{"id":"str","utterance":"str","context":"str",
"label_cat":["anger", "joy"],
"vad":{"v":0.00,"a":0.00,"d":0.00},
"vad_conf":0.0,
"qc_flags":{"len_ok":true,"tox":false,"example_unclear":false}}
A.5 Mixture Sampling (20:80 / 50:50 / 80:20).
python -u src/mix_sampler.py \
--goemo data/processed/goemotions.jsonl \
--empat data/processed/empathetic.jsonl \
--ratio 20:80 --vad_conf_min 0.80 --dev_frac 0.05 --seed 42 \
--outdir data/processed/mix_20_80
# Repeat for 50:50 and 80:20
A.6 Training Config (Qwen-1.8B-Chat).
# configs/sft_qwen1p8b.yaml (only essential fields)
base_model: /root/autodl-tmp/Emoloom-2B/models/qwen1_5_1_8b_chat/Qwen/Qwen1___5-1___8B-Chat
train_path: /root/autodl-tmp/Emoloom-2B/data/processed/mix_20_80/train.jsonl
dev_path: /root/autodl-tmp/Emoloom-2B/data/processed/mix_20_80/dev.jsonl
save_dir: /root/autodl-tmp/Emoloom-2B/outs/sft_qwen_mix2080
seed: 42
max_len: 1536
epochs: 1
per_device_train_batch_size: 1
gradient_accumulation_steps: 128
learning_rate: 1.2e-5
weight_decay: 0.05
warmup_ratio: 0.03
lr_scheduler_type: cosine
bf16: true
gradient_checkpointing: true
logging_steps: 10
save_steps: 800
save_total_limit: 2
report_to: none
Run:
python -u -m src.train_sft --cfg configs/sft_qwen1p8b.yaml | tee -a runs/sft_qwen_mix2080.log
A.7 Evaluation (Dev & Cross-Corpus).
# In-domain dev
python -u src/eval_dev.py \
--model_dir outs/sft_qwen_mix2080 \
--dev data/processed/mix_20_80/dev.jsonl \
--out outs/sft_qwen_mix2080_eval.json
# Cross-corpus (DailyDialog converted to /root/.../dev.jsonl)
python -u src/eval_dev.py \
--model_dir outs/sft_qwen_mix2080 \
--dev /root/autodl-tmp/Emoloom-2B/dev.jsonl \
--out outs/sft_qwen_mix2080_dd.json
A.8 Quick Eval with ETA (Time-Budgeted).
python -u src/eval_quick_eta.py \
--model_dir outs/sft_qwen_mix2080 \
--dev /root/autodl-tmp/Emoloom-2B/dev.jsonl \
--exp sft_qwen_mix2080_dd_quick \
--time_budget_min 60 --max_new_tokens 48 --ctx_max_chars 400
A.9 Ratio Comparison (Auto-Collect).
python -u src/compare_ratios.py \
--base_outs /root/autodl-tmp/Emoloom-2B/outs \
--exps sft_qwen_mix2080 sft_qwen_mix5050 sft_qwen_mix8020
A.10 Minimal One-Pass Checklist.
# Env -> Mix (x3) -> Train -> Eval(dev) -> Eval($D$ quick) -> Compare
All training uses use_cache=false and gradient checkpointing; OOM self-healing reduces max_len and increases grad_accum, then resumes.
Appendix B Appendix B: Extra Results
B.1 Qualitative Stability. Across seeds (three runs), dev curves show monotonic loss decay with early stabilization of gradient norms ($\sim$1.5–2.0). We observe fewer format failures as training proceeds, consistent with improved ParseOK in the main results.
B.2 Cross-Corpus Behavior. Under a one-hour budget on DailyDialog, the 20:80 model maintains usable Macro-F1 and a stable validity rate; variance primarily reflects domain mismatch (dialog style and topic shifts). Stronger valence robustness is observed under polarity flips, while arousal/dominance are less constrained by the augmentation (as intended).
B.3 Sensitivity (Qualitative). Removing the VAD-preserving loss tends to increase VAD RMSE; removing the appraisal verifier slightly degrades label consistency for fairness/controllability-related emotions; removing Valence Flip weakens polarity symmetry diagnostics. Cooler temperature schedules converge faster but may reduce coverage on minority emotions.
Appendix C Appendix C: Visualization Export Specs
C.1 General. Vector or high-resolution PNG; English labels; external legend; non-overlapping annotations; colorblind-safe palette.
C.2 Bars & Radar. Bars: grouped by mixture; error bars (std across seeds) when available. Radar: identical axis limits across mixtures; tick labels at uniform intervals; bold highlight for the best mixture.
C.3 Uncertainty Bands. For loss/metric curves: median line with 25–75% band; uniform smoothing window; no extrapolation beyond observed steps.
C.4 Reproducible Export. All plots exported with fixed DPI, font size, and bounding boxes; filenames referenced in the paper (e.g., dev_scores_bars.png, dev_scores_radar.png) are generated by the evaluation scripts.
Appendix D Appendix D: Ethics and Data Processing Details
D.1 Data Use & Privacy. We report aggregate metrics only; no raw text redistribution; IDs are anonymized. Cross-corpus evaluation is inference-only and respects original licenses.
D.2 QC & Filtering. We apply basic quality controls (length, language, toxicity flags) and require a minimum VAD confidence threshold ($\tau\in\{0.75,0.80\}$). Duplicates are removed via sha1(text||id).
D.3 Weak-Label Generation. Weak VAD for DailyDialog is computed by token-level NRC-VAD aggregation with missing tokens ignored and tail-trimmed normalization; negation and intensifiers are minimally handled via heuristic weight adjustments. Weak labels are used only for training-time regularization and diagnostics, not as gold labels.
D.4 Reproducibility Hygiene. Seeds for Python/NumPy/PyTorch are fixed; cudnn.deterministic=True, cudnn.benchmark=False. Checkpoints include config, tokenizer hash, and git commit for auditability. Metrics are reported on valid JSON outputs with ParseOK shown alongside to expose formatting failures.